### v2

@@ H.2 Analysis for Stochastic Mirror Descent
$\sum_{s=t_{1}}^{t_{2}}\eta_{s}\left[f(\mathbf{x}_{s})-f(\mathbf{x})\right]\leq B(\mathbf{x},\mathbf{x}_{t_{1}-1})-B(\mathbf{x},\mathbf{x}_{t_{2}})+\sum_{s=t_{1}-1}^{t_{2}-1}\left(\eta_{s+1}\langle\bm{\epsilon}_{s},\mathbf{x}-\mathbf{x}_{s}\rangle+\eta_{s+1}^{2}\sigma_{s}^{2}\right),$ (186)
@@
$q_{s}:=f(\mathbf{x}_{s})-f^{*},\ V_{s}:=\eta_{s+1}\langle\bm{\epsilon}_{s},\mathbf{x}_{*}-\mathbf{x}_{s}\rangle,\ U_{s}^{k}:=\eta_{s+1}\langle\bm{\epsilon}_{s},\mathbf{x}_{t-k}-\mathbf{x}_{s}\rangle$ (187)
for all $0\leq s\leq t$, where $\mathbf{x}_{*}\in\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x})$ (the existence of $\mathbf{x}_{*}$ can be implied by Assumptions 1 and 2). Obviously, we have $q_{s}\geq 0$ and $\left\{V_{s}\right\}_{s=0}^{t-1},\left\{U_{s}^{k}\right\}_{s=t-k+1}^{t-1}$ are martingale difference sequences. Then we have
@@
The above bound usually relates to the average-iterate. To incorporate Lemma 29, we consider the following relation:
@@
(${\leq}$ $B(\mathbf{x}_{t-k},\mathbf{x}_{t-k})-B(\mathbf{x}_{t-k},\mathbf{x}_{t})+\sum_{s=t-k}^{t-1}\left(\eta_{s+1}\langle\bm{\epsilon}_{s},\mathbf{x}_{t-k}-\mathbf{x}_{s}\rangle+\eta_{s+1}^{2}\sigma_{s}^{2}\right)$
@@
With these preparations, we invoke Lemma 29 to deduce
@@
Next, we utilize Lemma 28 to bound the sum of the martingale difference sequence. In view of $\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|\leq 2\sqrt{2}D,\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{X}$ (cf. (Yu et al., 2024, Proposition A.6)), we have
$V_{s}\leq\eta_{s+1}\left\|\bm{\epsilon}_{s}\right\|_{*}\left\|\mathbf{x}_{*}-\mathbf{x}_{s}\right\|\leq\frac{D}{\sigma_{s}\sqrt{s+1}}\sigma_{s}\cdot 2\sqrt{2}D=\frac{2\sqrt{2}D^{2}}{\sqrt{s+1}},$ (193)
$U_{s}^{k}\leq\eta_{s+1}\left\|\bm{\epsilon}_{s}\right\|_{*}\left\|\mathbf{x}_{t-k}-\mathbf{x}_{s}\right\|\leq\frac{D}{\sigma_{s}\sqrt{s+1}}\sigma_{s}\cdot 2\sqrt{2}D=\frac{2\sqrt{2}D^{2}}{\sqrt{s+1}}.$
where we make use of $\eta_{s}\leq D/(\sigma_{s-1}\sqrt{s})$. Hence, we invoke Lemma 28 to deduce that with probability at least $1-\delta/t$,
@@
(${\leq}$ $\frac{1}{t}\sum_{k=1}^{t-1}\frac{1}{k}\left(\sqrt{t}+\frac{1}{\sqrt{\ln t}}\right)\leq\frac{1+\ln t}{\sqrt{t}}+\frac{1+\ln t}{t\sqrt{\ln t}}$
@@
$F_{t}\leq$ $30\max\left\{1,\frac{\sigma_{s-1}}{G_{s-1}}\right\}L_{t-1}D^{2}\sqrt{\ln\frac{16}{\delta}}$ (202)
$+12D\sigma_{t-1}\sqrt{\ln\frac{16}{\delta}}+4D\sigma_{t-1}\sqrt{\ln\frac{2t^{3}}{\delta}}\ln t=\widetilde{F}_{t}.$
We will show $\widetilde{F}_{t}=\widetilde{O}(1)$ by induction in the sequel. For the base case, we have $G_{0}=\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}<\infty$. Hence $L_{0},\sigma_{0}$ are absolute constants. Since the high probability bound (201) holds, we clearly have $F_{1}\leq\widetilde{F}_{1}<\infty$ and $G_{1}\leq\widetilde{G}_{1}<\infty$ by Lemma 16. For the induction step, we have $G_{t-1}=\widetilde{O}(1)$ from the induction basis. Then under Assumptions 3 and 4,
@@
hold with probability at least $1-\delta$, which implies $F_{t}=\widetilde{O}(1)$. Then we invoke Lemma 16 again to deduce that $G_{t}=\widetilde{O}(1)$. Combining the base case and the induction step finishes the proof.

### v3

@@ Appendix I Analysis of Stochastic Mirror Descent
$\sum_{s=t_{1}}^{t_{2}}\eta_{s}\left[f(\mathbf{x}_{s})-f(\mathbf{x})\right]\leq B(\mathbf{x},\mathbf{x}_{t_{1}-1})-B(\mathbf{x},\mathbf{x}_{t_{2}})+\sum_{s=t_{1}-1}^{t_{2}-1}\left(\eta_{s+1}\langle\boldsymbol{\epsilon}_{s},\mathbf{x}-\mathbf{x}_{s}\rangle+\eta_{s+1}^{2}\sigma_{s}^{2}\right),$ (184)
@@
$q_{s}:=f(\mathbf{x}_{s})-f^{*},\ V_{s}:=\eta_{s+1}\langle\boldsymbol{\epsilon}_{s},\mathbf{x}_{*}-\mathbf{x}_{s}\rangle,\ U_{s}^{k}:=\eta_{s+1}\langle\boldsymbol{\epsilon}_{s},\mathbf{x}_{t-k}-\mathbf{x}_{s}\rangle$ (185)
for all $0\leq s\leq t$, where $\mathbf{x}_{*}\in\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x})$ (the existence of $\mathbf{x}_{*}$ can be implied by Assumptions 2.5 and J.2). Obviously, we have $q_{s}\geq 0$ and $\left\{V_{s}\right\}_{s=0}^{t-1},\left\{U_{s}^{k}\right\}_{s=t-k+1}^{t-1}$ are martingale difference sequences. Then we have
@@
The above bound usually relates to the average-iterate. To incorporate Lemma J.7, we consider the following relation:
@@
[…]
@@
With these preparations, we invoke Lemma J.7 to deduce
@@
Next, we utilize Lemma J.6 to bound the sum of the martingale difference sequence. In view of $\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|\leq 2\sqrt{2}D,\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{X}$ (cf. (Yu et al., 2024, Proposition A.6)), we have
$V_{s}\leq\eta_{s+1}\left\|\boldsymbol{\epsilon}_{s}\right\|_{*}\left\|\mathbf{x}_{*}-\mathbf{x}_{s}\right\|\leq\frac{D}{\sigma_{s}\sqrt{s+1}}\sigma_{s}\cdot 2\sqrt{2}D=\frac{2\sqrt{2}D^{2}}{\sqrt{s+1}},$ (191)
$U_{s}^{k}\leq\eta_{s+1}\left\|\boldsymbol{\epsilon}_{s}\right\|_{*}\left\|\mathbf{x}_{t-k}-\mathbf{x}_{s}\right\|\leq\frac{D}{\sigma_{s}\sqrt{s+1}}\sigma_{s}\cdot 2\sqrt{2}D=\frac{2\sqrt{2}D^{2}}{\sqrt{s+1}}.$
where we make use of $\eta_{s}\leq D/(\sigma_{s-1}\sqrt{s})$. Hence, we invoke Lemma J.6 to deduce that with probability at least $1-\delta/t$,
@@
$\texttt{term-A}=$ $\sum_{k=1}^{t-1}\frac{1}{kt}\cdot\frac{1}{\sqrt{\ln\left(\frac{t}{t-k}\right)}}+\sum_{k=1}^{t-1}\frac{1}{t(t-k)}\cdot\frac{1}{\sqrt{\ln\left(\frac{t}{t-k}\right)}}$ (196)
@@
[…]
We will show $\widetilde{F}_{t}=\widetilde{O}(1)$ by induction in the sequel. For the base case, we have $G_{0}=\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}<\infty$. Hence $L_{0},\sigma_{0}$ are absolute constants. Since the high probability bound (199) holds, we clearly have $F_{1}\leq\widetilde{F}_{1}<\infty$ and $G_{1}\leq\widetilde{G}_{1}<\infty$ by Lemma D.3. For the induction step, we have $G_{t-1}=\widetilde{O}(1)$ from the induction basis. Then under Assumptions 3.3 and J.1,
@@
hold with probability at least $1-\delta$, which implies $F_{t}=\widetilde{O}(1)$. Then we invoke Lemma D.3 again to deduce that $G_{t}=\widetilde{O}(1)$. Combining the base case and the induction step finishes the proof. ∎
Proof of Theorem J.3.
In the proof of Lemma J.9, we have shown that with probability at least $1-\delta$ (see (186), (189), (194) and (199)),
By Lemma J.9, we have $G_{t}=\widetilde{O}(1),\ L_{t}=\widetilde{O}(1),\ \sigma_{t}=\widetilde{O}(1)$ that holds with high probability along the trajectory. Therefore, we have $\sigma_{t-1}^{\max}=\widetilde{O}(1),\widetilde{L}_{t-1}^{\max}=\widetilde{O}(1)$. In view of $D_{t}=\widetilde{O}(1)$, we conclude our proof by combining these quantities:
Appendix K Analysis for Non-convex Composite Mirror Descent
First, we introduce a standard yet useful lemma.
Lemma K.1 (Lemma 6.4 in Lan (2020)).
Let $\mathcal{G}_{t}$ be defined in (16), it holds that
$\langle\nabla f(\mathbf{x}_{t}),\mathcal{G}_{t}\rangle\geq\left\|\mathcal{G}_{t}\right\|^{2}+\frac{\phi(\mathbf{x}_{t+1})-\phi(\mathbf{x}_{t})}{\eta}.$
The proof of Theorem 5.1 stems from the textbook analysis of mirror descent in the non-convex case (Lan, 2020, Theorem 6.5), together with our previous generalized smooth analysis.
Proof of Theorem 5.1.
First, we provide a quick and intuitive explanation for our proof outline. Recall Lemma E.2, which bounds the suboptimality gaps $\left\{f(\mathbf{x}_{t})-f^{*}\right\}$ along the optimization trajectory without convexity. This is exactly the key to the derivations in the non-convex case. Since we have $f(\mathbf{x}_{t})-f^{*}\leq F$ and $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G$ for all $t\in\mathbb{N}$, we immediately leverage the effective smoothness constant $L$ and reduce the analysis for $\ell*$-smooth functions to the textbook analysis of standard $L$-smooth functions. Below, we present formal reasoning by induction.
When $t=0$, the induction basis is trivially satisfied. Suppose for all $s\leq t-1$ we have bounded suboptimality gaps and bounded gradients. Now that the conditions of Lemma 2.7 is satisfied, we apply Lemma 2.7 to the $t$-th iterate:
$f(\mathbf{x}_{t+1})\leq$ $f(\mathbf{x}_{t})+\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}_{t}\rangle+\frac{L}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$
$\leq$ $f(\mathbf{x}_{t})-\eta\langle\nabla f(\mathbf{x}_{t}),\mathcal{G}_{t}\rangle+\frac{\eta^{2}L}{2}\left\|\mathcal{G}_{t}\right\|^{2}$
[…]
Since $\eta\leq 1/L$, after rearrangement we get
$f(\mathbf{x}_{t+1})\leq F(\mathbf{x}_{t+1})\leq F(\mathbf{x}_{t})-\frac{\eta}{2}\left\|\mathcal{G}_{t}\right\|^{2}\leq F(\mathbf{x}_{t})\leq\cdots\leq F(\mathbf{x}_{0}),$ (206)
which implies $f(\mathbf{x}_{t+1})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}+\phi(\mathbf{x}_{0})$. So we deduce $\left\|\nabla f(\mathbf{x}_{t+1})\right\|_{*}\leq G$ by Lemma D.3. Hence, throughout the optimization trajectory, the bounded suboptimality gaps as well as gradients hold uniformly. Thus, we can start from (206) to perform a telescoping:
$\frac{1}{T}\sum_{t=0}^{t-1}\left\|\mathcal{G}_{t}\right\|^{2}\leq\frac{F(\mathbf{x}_{0})-F(\mathbf{x}_{T})}{\eta T}\leq\frac{F(\mathbf{x}_{0})-F^{*}}{\eta T},$
which completes the proof. ∎